# Does SuperKamiokande Observe Lévy Flights of Solar Neutrinos?

### v1

@@ 3 Probability Density Function and Differential Equation: Lévy Flights
We consider a diffusion process generated by a waiting time pdf with a finite characteristic time $T$ that can be modeled with $a$ Poissonian distribution, and a jump length pdf $\lambda(x)$ given by a Lévy distribution with index $20<\alpha<$ 2 (Metzler and Klafter, 2000). The Fourier transform of $\lambda(x)$ is
@@
where ${}_{-\infty}D^{\alpha}_{x}$ is the fractional Weyl operator.
Upon Laplace inversion of Eq. (9), we get the characteristic function of the jump pdf
@@
where $c$ in the contour is such that $-\alpha<c<0$. Replacing $\frac{s}{\alpha}$ by $s$, observing that $\alpha$ coming from $s~{}{\rm d}s$ is canceled with $\alpha$ sitting outside, we have the following:
@@
×$\times\sum_{\nu=0}^{\infty}\frac{(-1)^{\nu}}{\nu!}\frac{1}{\Gamma(\frac{1+\alpha}{2}+\frac{\alpha}{2}\nu)\Gamma(1-\frac{\alpha}{2}(1+\nu))}\left(\frac{|x|^{\alpha}}{(2K)^{\alpha}t}\right)^{\nu},\alpha\neq 1,\alpha\neq 2.$
4 Discussion
The first solar neutrino experiment led by Raymond Davis Jr. showed a deficit of neutrinos relative to the solar model prediction, referred to as the solar neutrino problem since the 1970s. The Kamiokande experiment led by Masatoshi Koshiba successfully observed solar neutrinos, as first reported in 1980s. The solar neutrino problem was solved due to neutrino oscillations by comparing the Super-Kamiokande and Sudbury Neutrino Observatory results. While recent decades have offered tremendous advances in solar neutrinos across the fields of solar physics (Buldgen et al., 2024), nuclear physics (Bertulani et al., 2022), and neutrino physics (Slad, 2023), many lingering mysteries remain.
This paper takes advantage of publicly available solar neutrino Super-Kamiokande data and analyses them by applying diffusion entropy analysis and standard deviation analysis. The result is a scaling exponent $\delta<1$ indicating anomalous diffusion of solar neutrinos in terms of Lévy flights. Based on this result the paper developes the probability density function for neutrino flights and derives the respective differential equation in terms of a fractional Fokker-Planck equation. Accordingly, the closed form analytic representation of the neutrino power density function is given as a Fox H-function that can be used for further numerical exercises for the benefit of solar neutrino physics.
Acknowledgements
The authors are grateful for the support in the diffusion entropy analysis and standard deviation analysis by Dr. Alexander Haubold while he was doing his research at the Department of Computer Science, Columbia University, New York (USA). The results of diffusion entropy analysis were independently confirmed by the research of Dr. Nicy Sebastian, Department of Statistics, St Thomas College, Thrissur, University of Calicut, Kerala (India).
Figure 5: The Standard Deviation Analysis (SDA) of the $hep$ solar neutrino data from the SuperKamiokande I and II experiment.
Figure 6: The Diffusion Entropy Analysis (DEA) of the $hep$ solar neutrino data from the SuperKamiokande I and II experiment.

### v2

@@ 3.  Probability Density Function and Differential Equation: Lévy Flights
We consider a diffusion process generated by a waiting time pdf with a finite characteristic time $T$ that can be modeled with $a$ Poissonian distribution, and a jump length pdf $\lambda(x)$ given by a Lévy distribution with index $0<\alpha<2$ (Metzler and Klafter, 2000). The Fourier transform of $\lambda(x)$ is
@@
where $_{-\infty}D^{\alpha}_{x}$ is the fractional Weyl operator. Upon Laplace inversion of Eq. (9), we get the characteristic function of the jump pdf
@@
, where $c$ in the contour is such that $\alpha<c<0$. Replacing $\frac{s}{\alpha}$ by $s$, observing that $\alpha$ coming from $s~{\rm d}s$ is canceled with $\alpha$ sitting outside, we have the following:
@@
$\times\sum_{\nu=0}^{\infty}\frac{(-1)^{\nu}}{\nu!}\frac{1}{\Gamma(\frac{1+\alpha}{2}+\frac{\alpha}{2}\nu)\Gamma(1-\frac{\alpha}{2}(1+\nu))}\left(\frac{|x|^{\alpha}}{(2K)^{\alpha}t}\right)^{\nu},\alpha\neq 1,\alpha\neq 2.$
4.  Discussion
The first solar neutrino experiment led by Raymond Davis Jr. showed a deficit of neutrinos relative to the solar model prediction, referred to as the solar neutrino problem since the 1970s. The Kamiokande experiment led by Masatoshi Koshiba successfully observed solar neutrinos, as first reported in 1980s. The solar neutrino problem was solved due to neutrino oscillations by comparing the SuperKamiokande and Sudbury Neutrino Observatory results. While recent decades have offered tremendous advances in solar neutrinos across the fields of solar physics (Buldgen et al., 2024; Yang and Tian, 2024), nuclear physics (Bertulani et al., 2022; Hwang et al., 2023), and neutrino physics (Sturrock et al., 2021; Slad, 2024), many lingering mysteries remain.
This chapter takes advantage of publicly available solar neutrino SuperKamiokande data and analyses them by applying diffusion entropy analysis and standard deviation analysis. The result is a scaling exponent $\delta<1$ indicating anomalous diffusion of solar neutrinos in terms of Lévy flights. Based on this result the chapter developes the probability density function for neutrino flights and derives the respective differential equation in terms of a fractional Fokker-Planck equation. Accordingly, the closed form analytic representation of the neutrino power density function is given as a Fox H-function that can be used for further numerical exercises for the benefit of solar neutrino physics.
The authors have grateful for the support in the diffusion entropy analysis and standard deviation analysis by Dr. Alexander Haubold while he was doing his research at the Department of Computer Science, Columbia University, New York (USA). The results of diffusion entropy analysis were independently confirmed by the research of Dr. Nicy Sebastian, Department of Statistics, St Thomas College, Thrissur, University of Calicut, Kerala (India).